OpenStudy (australopithecus):

For the Power series \[\sum_{n=1}^{\infty} 3^{n}(x+4)^{n}/(n^{1/2})\] By applying the ratio test I get 3|x+4| so I determined a = -4 how do I find the radius (R) some how my prof got 1/3 I do not understand how he came upon such a solution

OpenStudy (turingtest):

the series will converge if \[3|x+4|<1\implies|x+4|=\frac13\]so R=1/3 is the radius of convergence

OpenStudy (australopithecus):

isn't the radius the distance of -4 to the end points of the interval?

OpenStudy (turingtest):

no, I think you are thinking of the interval of validity, which is\[-\frac13<x+4<\frac13\]\[-\frac{13}3<x<-\frac{11}3\]

OpenStudy (australopithecus):

man I'm confused lol

OpenStudy (australopithecus):

is there a general method to compute the radius?

OpenStudy (turingtest):

look at what the ratio test is looking for...

OpenStudy (australopithecus):

it is looking for the area of convergence, for which values of x make the function convergent (although the end points need to be checked)

OpenStudy (australopithecus):

sorry by function I mean series

OpenStudy (turingtest):

when\[\lim_{n\to\infty}\left|{\{a_{n+1}\}\over\{a_n\}}\right|<1\]we get convergence, so we take that limit, and if the limit exists we call it \(L\) we then have something that looks like\[L|f(x)|<1\]the radius of convergence is the range in which \(|f(x)|\) coonverges, which is\[|f(x)|<\frac1L\]

OpenStudy (turingtest):

...to find the interval of validity you would solve the above for x

OpenStudy (australopithecus):

ok thanks :) I think I get it now

OpenStudy (turingtest):

just in case you want further info http://tutorial.math.lamar.edu/Classes/CalcII/PowerSeries.aspx

OpenStudy (australopithecus):

thanks for the link I will definitely check it out

OpenStudy (turingtest):

welcome!